calculate the mass of chlorine gas contained in a 5.00dm^3 flask at 25°C and 98 kPa pressure
1
Expert's answer
2018-03-09T10:01:07-0500
According to ideal gas law:
m/M×RT=PV
Find the mass:
m = (M×P×V)/(R×T)
V = 0.005 m3
T = 298 K
P = 98000 Pa
R = 8.314 J×mol-1×K-1
M = 71 g×mol-1
m = (71×98000×0.005)/(298×8.314) = 14.04 (g)